# OpenFly: A Versatile Toolchain and Large-scale Benchmark for Aerial Vision-Language Navigation

### v4

@@ V-C2 Visual Token Merging
For each set of candidate keyframes obtained in the previous selection process, a visual encoder maps each input image to multiple visual tokens, with each token representing the information of an image patch. Considering the potential inter-frame patch redundancy, we take a strategy that similar tokens in subsequent adjacent frames are periodically merged. Specifically, we select the first frame in a keyframe set as the reference, since it usually contains the crucial observation indicating the time for action transition. Then, we densely calculate the cosine similarities between each pair of visual tokens of the reference image and the subsequent image. Next, we merge the tokens with high similarity by averaging them. The unmerged tokens in the subsequent frame will be discarded. The merging operation is iteratively performed until the entire keyframe set has been traversed. Besides, we maintain a memory bank with a capacity of $K$ images, which follows a first-in-first-out (FIFO) policy to retain the latest keyframes.
After the above process, $M$ visual tokens $E=\{e_{1},e_{2},\cdots,e_{M}\}$ are obtained for each set of keyframes. Since aerial VLN requires UAVs to perform long-distance flights based on instructions, we continue to carry out token compress to reduce the computational burden. The compressed visual tokens $E_{c}$ are obtained through grid pooling [27]. Notably, we keep the visual tokens of the current frame uncompressed to capture the latest visual observation, as it contains the most important information for flight action prediction.
V-C3 Action Prediction
[…]
TABLE IV: Comparison results on the test-unseen split.
[…]
TABLE V: Ablation study on the test-seen split. ‘KS’ and ‘VTM’ denote keyframe selection and visual token merging, respectively.
[…]
VI Experiments
VI-A Implementation and Training Details
The proposed OpenFly-Agent is composed of the Dino-SigCLIP (224 $\times$ 224 pixels) as a vision encoder and the pretrained Llama-2 (7B) as a language model. The visual tokens extracted by Dino-SigLIP are aligned to the same dimension as text embeddings using a projection layer, and then fed into the Llama-2 model. The current frame during flight remains 256 tokens and all historical keyframes are compressed into 1 token. The capacity $K$ of the history memory bank is set to 2 in our experiment. For action prediction, the last 256 tokens in the vocabulary are used as special tokens for action representation.
VI-B Evaluation Metrics
Four standard metrics in VLN tasks are adopted to evaluate different methods, i.e., navigation error (NE), success rate (SR), oracle success rate (OSR), and success weighted by path length (SPL). NE measures the average deviation between the UAV’s final stopping point and the ground-truth destination. SR calculates the proportion of successful tasks, where a task is considered successful if the UAV stops within 20 m of the target. In OSR, if any point on the trajectory is within 20 m of the target, the task can be considered successful. SPL calculates the success rate weighted by the ratio of the ground-truth path length to the actually-executed path length.
VI-C Results
VI-C1 Quantitative results
We test the proposed OpenFly-Agent on the test-seen and test-unseen splits and differentiate among easy, moderate, and hard difficulty levels. Multiple VLN methods are compared with ours, where ‘Random’ means an agent randomly selects one action to execute until the ‘stop’ action is chosen, Seq2Seq [23] adopts a simple recurrent model for action sequence prediction, CMA [23] uses a bi-directional LSTM to implement cross-modal attention between instructions and images, AerialVLN [29] combines CMA and a look-ahead guidance strategy to enhance robustness. Navid [56] leverages the strong ability of VLMs and develops a video-based architecture, achieving impressive results in indoor VLN.
Table. III shows quantitative results on the test-seen split. Seq2Seq, CMA, and AerialVLN achieve limited success rates. Counterintuitively, they perform worse on easy samples. We find that their OSR is not low while they are difficult to stop at a short-range distance. In contrast, Navid obtains better results, demonstrating the great potential of VLMs in aerial VLN. Our OpenFly-Agent achieves the best performance for its strong VLM backbone and the proposed strategies, while there is still much room for improvement. Table. IV presents quantitative results on the test-unseen split, illustrating the generalization capabilities of these VLN methods. Similarly, our method achieves better performance, exhibiting a certain degree of robustness. However, all methods are significantly degraded, which indicates models with greater generalization are urgently needed to be developed.
@@ VI-C1 Quantitative results
Figure 7: Illustration of an aerial VLN trajectory generated by OpenFly-Agent, which successfully predicts actions following the instruction when encountering landmarks.
@@
VI-C2 Qualitative results
Fig. 7 presents a qualitative result, where our OpenFly-Agent successfully navigates to the destination according to the instruction. It presents a powerful capability in perceiving environments and aligning observations with complex instructions. In addition, Fig. 8 shows two failure cases, where our model fails to identify the landmark or output actions with proper amplitudes.
VI-D Ablation Study
Ablation studies are conducted to evaluate the contribution of the keyframe selection and visual token merging in OpenFly-Agent. Table. V shows the results, where OpenVLA [22] is our baseline. Using only the current frame as an observation makes OpenVLA perform poorly in the aerial VLN task. From ‘History + VTM’ we can see that historical frames significantly improve the success rate. The keyframe selection strategy further increases the SR from 12.7% to 18.5%, demonstrating the effectiveness of key observations. Besides, the comparison between ‘KS’ and ‘KS + VTM’ indicates the great effect of our visual token merging strategy.
VII Limitations and Future Work
The proposed OpenFly platform incorporates various rendering engines/techniques to provide high-quality scenes. Specifically, this is the first attempt to use 3D GS reconstructed scenes to support real-to-sim training and testing, while in the reconstruction of large-scale areas, a few visual artifacts are inevitably present. Future work will focus on exploring more effective reconstruction methods to enhance realism in large-scale scenes. Besides, the proposed OpenFly-Agent is built upon the large VLN model architecture, which is not practical for real-time deployment on UAVs. To address this, future research should focus on developing more efficient architectures and effective quantization techniques.
VIII Conclusion
In this work, we present OpenFly, a platform designed for large-scale data collection in aerial Vision-and-Language Navigation (VLN). OpenFly integrates multiple rendering engines and advanced real-to-sim techniques for data generation, enabling efficient collection of diverse, high-quality aerial VLN data. The resulting large-scale dataset comprises 100k trajectories across 18 distinct scenes, spanning a wide range of altitudes and difficulty levels, which is significantly superior than existing ones. Furthermore, we propose OpenFly-Agent, a keyframe-aware aerial navigation model capable of directly predicting flight actions based on observations and language instructions. Extensive experiments validate the effectiveness of the proposed method, and establishing a comprehensive benchmark for future advancements in aerial navigation.

### v5

@@ Appendix I Implementation and Training Details
Appendix J Qualitative Experimental Results
Fig. 8 presents a qualitative result in a UE scene, where our OpenFly-Agent successfully navigates to the destination according to the instruction. It presents a powerful capability in perceiving environments and aligning observations with complex instructions. Fig. 9 presents another successful aerial VLN example in a 3D GS scene. The image style, flight heights, and viewpoints are significantly different from UE’s scenarios. In this case, our OpenFly-Agent exhibits robustness to handle data with great diversity. In addition, Fig. 10 shows two failure cases, where our model sometimes fails to identify the landmark or output actions with proper amplitudes.
@@
Figure 8: Illustration of an aerial VLN trajectory generated by OpenFly-Agent in a UE scene, which successfully predicts actions following the instruction when encountering landmarks.
Table 6: Results of real-world experiments.
model SR $\uparrow$ OSR $\uparrow$
Navid [16] 13.0% 17.39%
NaVILA [43] 13.0% 21.74%
OpenFly-Agent (Ours) 26.09% 34.78%
Table 7: Performance of OpenFly-Agent trained on different datasets.
dataset SR $\uparrow$ OSR $\uparrow$
AerialVLN  [17] 4.3% 13.04%
OpenFly (Ours) 26.09% 34.78%
Figure 9: Illustration of aerial VLN trajectories generated by OpenFly-Agent in a 3D GS scene.
@@
Appendix K Real-world Experiments
The real-world experiments are conducted in 23 real outdoor scenes, where each scene corresponds to an unseen VLN task created by human operators, and the trajectory lengths range from 50m to 500m. We use a Q250 airframe as a real agent, carrying an NVIDIA Jetson Xavier NX running Ubuntu 18.04 as the onboard computer. All methods run on an external PC communicating with the onboard computer to transfer images and action instructions. Two most recent models, Navid [16] and NaVila [43], are evaluated for comparison. The results are shown in Table 6, where our model achieves the best performance with 26.09% SR and 34.78% OSR, significantly outperforming the comparison methods. This experiment again indicates the superiority of our OpenFly-Agent. Besides, we also trained our model on both our own dataset and the AerialVLN dataset separately. The results are shown in Table 7, where the first row is trained on the AerialVLN dataset (without real-to-sim data), and the second row is trained on our OpenFly dataset. It strongly demonstrates the capability of our data generation method in bridging the sim-to-real gap. A qualitative result is presented in Fig. 11, and a dynamic demo can be found in our supplementary video.
Figure 11: Snapshots of the real-world experiment.